OpenStudy (anonymous):

Monday: Which is equivalent to (i+3)+i(2i−4)? a. −5i+1 b. −i+3 c. i−3 −3i+1

rvc (rvc):

remember : \(\rm\huge i^2=-1 \)

OpenStudy (michele_laino):

here we can apply the distributive property of multiplication over addition, so we can write this: \[\Large \left( {i + 3} \right) + i\left( {2i - 4} \right) = i + 3 + \left( {i \cdot 2i} \right) - 4i = ...?\]

OpenStudy (michele_laino):

also, please keep in mind the reply of @rvc